OpenStudy (anonymous):

what is the value of 10^5 in standord form ?

OpenStudy (turingtest):

10^1=10 10^2=10*10=100 10^3=10*10*10=? notice a pattern?

OpenStudy (anonymous):

=1000

OpenStudy (turingtest):

good, so we have 10^1=10 10^2=100 10^3=1000 do you see a pattern between the number of zeros in standard form and the exponent on the 10 ?

OpenStudy (anonymous):

Its 100,000 as my answer thank you :)

OpenStudy (turingtest):

welcome :) in general, 10^n can be written as a 1 followed by n zeros. Good for future maths ;)

OpenStudy (turingtest):

oh and, that's not the answer :P

OpenStudy (anonymous):

-.-

OpenStudy (turingtest):

look at the pattern again 10^1=10 10^2=100 10^3=1000 10^4=?

OpenStudy (anonymous):

10,000

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

and then next is 100,000

OpenStudy (turingtest):

oh I;m sorry, misread your question :P my bad, had it right the first time, sorry :(

OpenStudy (anonymous):

Lol its ok :)